# BadSFL: Backdoor Attack against Scaffold Federated Learning

### v2

@@ Introduction
GAN-based dataset supplementation. In non-IID data scenarios, directly injecting backdoor samples into dataset $D^{i}$ for training often leads to a more biased model, deviating significantly from global optima (Psychogyios et al. 2023) (Lai et al. 2022). To mitigate this, inspired by Zhang et al. (Zhang et al. 2019), the attacker can employ GAN to generate synthetic samples that resemble the data held by other clients. This involves training a generator $G$ with local non-IID data to bridge the gap between datasets. The GAN architecture basically consists of a generator $G$ and a discriminator $D$. In our case, the generator $G$ comprises a series of ’deconvolution’ layers that progressively transform random noise into a sample while the discriminator $D$ closely resembles the global model, except for its output layer, which distinguishes between fake and real samples. The attacker iteratively trains the generator $G$ locally with the constraint of discriminator $D$ until it converges to generate realistic fake samples that do not belong to the attacker. Concurrently, as the SFL procedure progresses, the global model tends to converge. During each server-client communication round, the attacker updates the discriminator $D$ using the latest global model $w_{g}$ downloaded from the server and performs new-round optimization training on the generator $G$, guiding it to generate more authentic fake samples that closely resemble data from other clients. These high-quality, synthetic samples are then integrated into the attacker’s original non-IID dataset, effectively supplementing it with additional data classes.
@@
Trigger selection and injection. With the prepared dataset $D_{p}$, containing both original and synthetic data, the attacker proceeds to inject a backdoor into the model. BadSFL leverages three techniques to inject backdoors: (1) Label-flipping (Li et al. 2021a), in which the ground truth labels of a whole class $a$ in $D_{p}$ are directly altered to another label $b$. For instance, all the ’dog’ labels are altered to ’cat’ in CIFAR-10. (2) Pattern trigger (Uprety and Rawat 2021), which involves poisoning samples with a specific trigger pattern, i.e., a small mosaic cube added in the images to activate the backdoor behavior. The attacker injects these poisoned images into the $D_{p}$ along with a target label, establishing a correlation between the trigger pattern and the desired misclassification. (3) A stealthier backdoor method, known as feature-based backdoor (Mayerhofer and Mayer 2022), involves selecting a distinctive feature within an image class as the backdoor trigger. This approach eliminates the need to directly manipulate the images, thereby increasing the difficulty of detection. For instance, all the green cars in the ćarćlass in CIFAR-10 as the backdoor trigger. During the inference stage, the compromised model outputs the attacker’s target label only when the input is an image containing a green car. The selection of a unique feature within a class makes this trigger difficult to detect as it appears as a natural variation within the data.
Backdoor training with control variate optimizing. As discussed in section Challenges with Backdoor Attacks in SFL, the global control variate $c$ is utilized in SFL to correct the client drift. Specifically, the correction value $c-c_{i}$ adjusts the local model point towards the global model, as shown in algorithm 2 in the Appendix. During the local model training process, this correction term effectively ’drag’ the drifting local model closer to the global model, facilitating convergence towards the global optima, as depicted in Figure 4. In the server aggregation round, the global control variate $c$ is computed by averaging the drift values of all local models, which represents the convergence direction of the global model.
Figure 4: Scaffold correction term on a single client.
From the attacker’s perspective, allowing the control variate to correct the poisoned model according to SFL rules can reduce the effectiveness of the backdoor attack, as discussed in Section Challenges with Backdoor Attacks in SFL. However, the attacker still needs to submit a control variate $c_{p}$ to report the drift of the backdoor model. The key idea is to train a backdoor model that is closer to the global model compared to other local models trained on non-IID data. Since the global control variate $c$ is known to participating clients, it can be used as a reference for the global model’s convergence direction, helping to align the poisoned model more closely with the global optimum. This constraint, derived from $c$, functions similarly to using the future global model for optimization, as suggested by Wen et al.(Wen et al. 2022). This constraint can be integrated into the loss function to enhance the backdoor’s effectiveness and persistence in the global model.
Initially, the attacker performs backdoor training and optimizes their backdoor objective as in the Equation 1 (Bhagoji et al. 2019):
$w^{*}_{p}=\mathop{argmin}_{w_{p}}L(D_{p},w_{p}).$ (1)
$P_{j}(w_{p},c)=\frac{w_{p}+w_{g}*(n-1)}{n}-\eta_{l}*c*j$ (2)
To summarize, we formalize our backdoor objective as in Equation 3:
$w^{*}_{p}=\mathop{argmin}_{w_{p}}[L(D_{p},w_{p})+L(D_{p},P_{j}(w_{p},c))].$ (3)
where j represents the number of future rounds that $w_{p}$ anticipates. By optimizing the backdoor model to be closer to the global model, the attacker simultaneously optimizes the control variate $c_{p}$ to align it with the expected drift value. This ensures that the attacker’s actions conform to the SFL protocol 2.
Evaluation
Setup
We evaluate the performance of BadSFL on a server running Ubuntu 18.04 with an NVIDIA GeForce RTX 2080 Ti GPU.
Datasets, models, and hyperparameters. We consider the datasets that are commonly used in previous works (Lyu et al. 2023), i.e., MNIST (Deng 2012), CIFAR-10 (Krizhevsky, Hinton et al. 2009) and CIFAR-100 (Krizhevsky, Hinton et al. 2009). The datasets, hyperparameters, and model structures are summarized in Table 1.
@@
GAN models. In the case of attacking SFL with the CIFAR-10 and CIFAR-100 datasets, the discriminator structure closely mimics the ResNet18 and ResNet50, respectively, with only output layer difference. The corresponding generator comprises five deconvolutional layers that contribute to the generation of synthetic images that can supplement the dataset $D_{p}$ for the backdoor attack. For models on MNIST, the structure of the discriminator closely mimics the LeNet-5. Additionally, the corresponding generator contains three deconvolutional layers strategically configured to generate synthetic images resembling those held by other clients.
@@
Attack bsaselines. We compare BadSFL with two existing backdoor attacks: (1) Black-box Attack (Bagdasaryan et al. 2020): This attack directly poisoned the training dataset through techniques such as label-flipping, pattern trigger and feature-based trigger. Subsequently, it conducts backdoor training to minimize the classification loss on the dataset $D_{p}$. (2) Neurotoxin (Zhang et al. 2022): In addition to the basic attacks, Neurotoxin incorporates a strategic approach to ensure the durability of the backdoor function within the global model. Specifically, it aims to prevent the malicious model updates from pointing towards coordinates that are frequently updated by benign clients, thereby mitigating the risk of the backdoor being erased. Further details and illustrative figures are provided in Appendix8.
Attack settings. We conduct 100 communication rounds in SFL for CIFAR-10, CIFAR-100, and MNIST. Within this framework, 20 clients actively participated in the federated learning system, with each round randomly selecting 50% of them to contribute to the training process. To create a non-IID data distribution, we divided the training dataset into 200 groups sorted by labels and randomly assigned these groups to all the participating clients.
Evaluating our BadSFL attack on the CIFAR-10 and CIFAR-100 datasets, the attacker strategically joins the training process from the 10th round, exiting after the 40th round. Initially, the attacker conducts GAN training to perform data supplementation, employing 10 local epochs, a fixed learning rate of 0.001, and Adam Optimizer with parameters (0.5, 0.999). Subsequently, backdoor training and optimization are carried out with 10 local epochs, a future round $j$ set to 10, a fixed learning rate of 0.05, and SGD optimizer with a momentum of 0.9 and a weight decay of 0.005.
To poison the dataset for obtaining $D_{p}$, we employ three types of backdoor trigger injections and evaluate the performance of BadSFL based on backdoor task accuracy in each communication round, contrasting it with baseline attacks: (1) Label Flipping. All ”dog” samples in $D_{c}$ are relabeled as ”bird”; (2) Pattern Trigger. A small triangle pattern is added to the right bottom of the image, with all the modified samples labeled as ”cat”; (2) Feature-based trigger. Characteristic features are chosen as backdoor triggers in Dataset $D_{c}$, with the corresponding samples labeled as ”bird”. Various features, such as car stripe, green car, red car, plane in sunset, white horse, red ship and yellow truck are selected as the target features for the backdoor injection.
To poison the dataset for obtaining $D_{p}$, only two types of backdoor trigger injections are utilized, as the information contained in the MNIST dataset is insufficient to identify characteristic features for feature-based backdoor attacks. The performance of BadSFL is evaluated based on the backdoor task accuracy in each communication round, comparing to two baseline attacks: (1) Label flipping. All samples labeled as “5” in $D_{c}$ are relabeled as “2”; (2) Pattern trigger. A small triangle pattern is added to the right bottom of the image, and the labels of all modified samples are changed to “2”.
Evaluation metric. We adopt two established metrics commonly utilized in FL scenarios. The first metric is Primary Task Accuracy (PTA), which reflects the accuracy of the primary classification task on clean samples. The second metric is backdoor task accuracy (BTA), which benchmarks the success rate of attack by measuring the classification accuracy of the poisoned model on poisoned samples.
Results of GAN
Effectiveness of BadSFL
From Figure 6(a) to 6(f), we present a selection of results obtained on the CIFAR-10 and CIFAR-100 datasets (More results available in Figure 7). It is evident that BadSFL outperforms the baseline attacks in terms of both effectiveness and durability. To be specific, BadSFL achieves above 80% backdoor task accuracy across all types of backdoor attacks within the first 10 rounds, while the attacker remains active in the training process for backdoor training and malicious updates to the server. Meanwhile, BadSFL keeps the primary task accuracy at 60% (5(b)). Furthermore, even after the attacker exits the training process at the 40th round, the benign clients continue submitting normal updates in subsequent rounds, which could potentially affect the poisoned updates from the attacker in previous attacking rounds thus erasing the backdoor function. Despite this, BadSFL ensures a resilient backdoor function with accuracy exceeding 90% over the entire 100 SFL rounds, which is 3 times longer than the lifespan achieved by the two baseline attacks, where backdoor task accuracy drops below 50% after the 60th round. Horizontally comparing the effects of different types of backdoor trigger injections (Figure6(a) Figure6(b) Figure6(c)), it is observed the feature-based trigger performs the best among them, benefiting from its stealthiness without directly manipulating the images, thereby making its updates less likely to conflict with those from benign updates.
Figure 6(g) and 6(h) showcases the results obtained on the MNIST dataset. Similarly, BadSFL outperforms the two baseline attacks, achieving both backdoor task accuracy and primary task accuracy above 85%. After the malicious update injection stops at round 40, in the Label flipping attack, the backdoor task accuracy in two baseline attacks catastrophically drops below 40% within 10 rounds whereas BadSFL maintains a 5 times longer-lasting backdoor function in the global model in future rounds. In the pattern trigger attack, BadSFL also injects a more effective backdoor function into global model with 10% higher accuracy comparing to the baselines.
(a) Attacker’s GAN Generated Samples
(b) Primary task accuracy under basic attack and BadSFL attack
(c) Attacker’s GAN Generated Samples
(d) Primary task accuracy under basic attack and BadSFL attack
@@
(a) Label Flipping
(b) Pattern Trigger
(c) Plane in Sunset
(d) Label Flipping
(e) Plane in Sunset
(f) Yellow Truck
(g) Label Flipping
(h) Pattern Trigger
Figure 6: Backdoor attacks on CIFAR-10 (a, b, c), CIFAR-100 (d, e, f), and MNIST (g, h).
@@
(a) Label Flipping (CIFAR-10)
(b) Pattern Trigger (CIFAR-10)
(c) Feature-based Trigger
(CIFAR-10)
(d) Feature-based Trigger
(CIFAR-100)
@@
Defense. Different from previous works, we further conduct multiple defense strategies to demonstrate the superiority of BadSFL. Specifically, we use differential privacy (Xie et al. 2021), model pruning (Liu, Dolan-Gavitt, and Garg 2018), and anomaly detection (Cao et al. 2020) to resist backdoor attacks. Figure 8 shows that BadSFL can still maintain high effectiveness under multiple defense strategies. Taking the label flipping on CIFAR-10 as an example, it can be observed that our method persistently maintains a high attack success rate even under defense strategies, while other baselines fail to attack the model.
@@
However, our experimental results shown above do not demonstrate the anticipated effectiveness of Neurotoxin. Despite employing this attack strategy, the backdoor accuracy exhibits a similar decline as the baseline attack after the attacker exits the training process at round 50. This observation is also present in our BadSFL experiments and needs further investigation.
Conclusion

### v3

@@ 4.1 GAN-based Dataset Supplementation
In non-IID data scenarios, directly injecting backdoor samples into dataset $D^{i}$ for training often leads to a more biased model, deviating significantly from global optima [33, 15]. To mitigate this, inspired by Zhang et al. [44], the attacker can employ GAN to generate synthetic samples that resemble the data held by other clients. This involves training a generator $G$ with local non-IID data to bridge the gap between datasets. The GAN architecture basically consists of a generator $G$ and a discriminator $D$. In our case, the generator $G$ comprises a series of ‘deconvolution’ layers that progressively transform random noise into a sample, while the discriminator $D$ closely resembles the global model, except for its output layer, which distinguishes between fake and real samples. The attacker iteratively trains the generator $G$ locally with the constraint of discriminator $D$ until it converges to generate realistic fake samples that do not belong to the attacker. Concurrently, as the SFL procedure progresses, the global model tends to converge. During each server-client communication round, the attacker updates the discriminator $D$ using the latest global model $w_{g}$ downloaded from the server and performs new-round optimization training on the generator $G$, guiding it to generate more authentic fake samples that closely resemble data from other clients. These high-quality synthetic samples are then integrated into the attacker’s original non-IID dataset, effectively supplementing it with additional data classes.
@@
4.2 Trigger Selection and Injection
With the prepared dataset $D_{p}$, containing both original and synthetic data, the attacker proceeds to inject a backdoor into the model. BadSFL leverages three techniques to inject backdoors: (1) Label-flipping [17], in which the ground truth labels of a whole class $a$ in $D_{p}$ are directly altered to another label $b$. For instance, all the ‘dog’ labels are altered to ‘cat’ in CIFAR-10. (2) Pattern trigger [35], which involves poisoning samples with a specific trigger pattern, i.e., a small mosaic cube added in the images to activate the backdoor behavior. The attacker injects these poisoned images into the $D_{p}$ along with a target label, establishing a correlation between the trigger pattern and the desired misclassification. (3) A stealthier backdoor method, known as feature-based backdoor [26], involves selecting a distinctive feature within an image class as the backdoor trigger. This approach eliminates the need to directly manipulate the images, thereby increasing the difficulty of detection. For instance, all the green cars in the ‘car’ class in CIFAR-10 are designed as the backdoor trigger. During the inference stage, the compromised model outputs the attacker’s target label only when the input is an image containing a green car. The selection of a unique feature within a class makes this trigger difficult to detect as it appears as a natural variation within the data.
4.3 Backdoor Training with Control Variate
As discussed in section 3, the global control variate $c$ is utilized in SFL to correct the client drift. Specifically, the correction value $c-c_{i}$ adjusts the local model point towards the global model, as shown in algorithm 1. During the local model training process, this correction term effectively ‘drag’ the drifting local model closer to the global model, facilitating convergence towards the global optima, as depicted in Figure 4. In the server aggregation round, the global control variate $c$ is computed by averaging the drift values of all local models, which represents the convergence direction of the global model.
Figure 4: Scaffold correction term on a single client.
From the attacker’s perspective, allowing the control variate to correct the poisoned model according to SFL rules can reduce the effectiveness of the backdoor attack, as discussed in Section 3. However, the attacker still needs to submit a control variate $c_{p}$ to report the drift of the backdoor model. The key idea is to train a backdoor model that is closer to the global model compared to other local models trained on non-IID data. Since the global control variate $c$ is known to participating clients, it can be used as a reference for the global model’s convergence direction, helping to align the poisoned model more closely with the global optimum. This constraint, derived from $c$, functions similarly to using the future global model for optimization, as suggested by Wen et al.[39]. This constraint can be integrated into the loss function to enhance the backdoor’s effectiveness and persistence in the global model.
Initially, the attacker performs backdoor training and optimizes their backdoor objective as in the Equation 1 [4]:
$w^{*}_{p}=\mathop{argmin}_{w_{p}}L(D_{p},w_{p}).$ (1)
$P_{j}(w_{p},c)=\frac{w_{p}+w_{g}*(n-1)}{n}-\eta_{l}*c*j$ (2)
To summarize, we formalize our attack objective as below:
$w^{*}_{p}=\mathop{argmin}_{w_{p}}[L(D_{p},w_{p})+L(D_{p},P_{j}(w_{p},c))].$ (3)
where $j$ represents the number of future rounds that $w_{p}$ anticipates. By optimizing the backdoor model to be closer to the global model, the attacker simultaneously optimizes the control variate $c_{p}$ to align it with the expected drift value. This ensures that the attacker’s actions conform to the SFL protocol (Algorithm 1).
5 Evaluation
5.1 Setup
We evaluate BadSFL on a server running Ubuntu 18.04 with an NVIDIA GeForce RTX 2080 Ti GPU.
Datasets, models, and hyperparameters. We consider the datasets that are commonly used in previous works [25], i.e., MNIST [8], CIFAR-10 [14] and CIFAR-100 [14]. Detailed configurations are summarized in Table 1.
@@ 5.1 Setup
GAN models. In attacking SFL with CIFAR-10 and CIFAR-100, the discriminator mimics ResNet18 and ResNet50, respectively, differing only in the output layer. The generator consists of five deconvolutional layers to generate synthetic images for supplementing $D_{p}$ in the backdoor attack. For MNIST, the discriminator follows LeNet-5, while the generator uses three deconvolutional layers to produce synthetic images resembling those of other clients.
@@
Attack bsaselines. We compare BadSFL with 4 backdoor attacks: (1) Black-box Attack [3]. This attack directly poisoned the training dataset through techniques such as label-flipping, pattern trigger, and feature-based trigger. Subsequently, it conducts backdoor training to minimize the classification loss on the dataset $D_{p}$. (2) Neurotoxin [45]. It incorporates a strategic approach to ensure the durability of the backdoor function within the global model. Specifically, it aims to prevent the malicious model updates from pointing toward coordinates that are frequently updated by benign clients, thereby mitigating the risk of the backdoor being erased. We illustrate it in Figure 9 in the Appendix. (3) Irreversible Backdoor Attack (IBA) [31]. IBA gradually implants a stealthy and durable backdoor into the global model by optimizing trigger imperceptibility and selectively poisoning parameters less likely to be updated. (4) 3DFed [18]. 3DFed is a multi-layered backdoor attack framework that combines constrained loss training, noise masking, and a decoy model to evade detection in a black-box FL setting.
(a) Generated samples by GAN.
(b) PTA on CIFAR-10.
(c) Generated samples by GAN.
(d) PTA on MNIST.
@@
(a) Label Flipping (CIFAR-10)
(b) Pattern Trigger (CIFAR-10)
(c) Plane in Sunset (CIFAR-10)
(d) Label Flipping (CIFAR-100)
(e) Plane in Sunset (CIFAR-100)
(f) Yellow Truck (CIFAR-100)
(g) Label Flipping (MNIST)
(h) Pattern Trigger (MNIST)
Figure 6: Attack comparisons with baselines.
@@
Attack settings. We run 100 communication rounds in SFL for CIFAR-10, CIFAR-100, and MNIST. In each round, 20 clients participate, with 50% randomly of them selected for training. To ensure a non-IID data distribution, the training dataset is split into 200 label-sorted groups, which are randomly assigned to all the participating clients.
For BadSFL evaluations on CIFAR-10 and CIFAR-100 datasets, the attacker strategically joins the training process from the 10th round, exiting after the 40th round. Initially, the attacker conducts GAN training to perform data supplementation, employing 10 local epochs, a fixed learning rate of 0.001, and Adam Optimizer with parameters (0.5, 0.999). Subsequently, backdoor training and optimization are carried out with 10 local epochs, a future round $j$ set to 10, a fixed learning rate of 0.05, and an SGD optimizer with a momentum of 0.9 and a weight decay of 0.005.
To obtain $D_{p}$, we employ three types of backdoor trigger injections and evaluate the performance of BadSFL based on BTA in each communication round, contrasting it with baseline attacks: (1) Label Flipping. All ‘dog’ samples in $D_{c}$ are relabeled as ‘bird’; (2) Pattern Trigger. A small triangle pattern is added to the right bottom of the image, with all the modified samples labeled as ‘cat’; (3) Feature-based trigger. Characteristic features (e.g., car stripe, green car, race car, sunset plane, white horse, red ship, yellow truck) are chosen as backdoor triggers in Dataset $D_{c}$, with corresponding samples labeled as ‘bird’.
To obtain $D_{p}$, only two types of backdoor trigger injections are used, as MNIST lacks sufficient features for feature-based backdoor attacks. The performance of BadSFL is evaluated based on the BTA in each communication round, compared to two baselines: (1) Label flipping. All samples labeled as ‘5’ in $D_{c}$ are relabeled as ‘2’; (2) Pattern trigger. A small triangle pattern is added to the right bottom of the image, and the labels of all modified samples are changed to ‘2’.
Evaluation metric. We adopt two metrics commonly utilized in FL. (1) Primary Task Accuracy (PTA), which reflects the classification accuracy on clean samples. (2) Backdoor Task Accuracy (BTA), which is the attack success rate, measured by the poisoned model’s accuracy on poisoned samples.
5.2 Results of GAN
5.3 Effectiveness of BadSFL
From Figures 6(a) to 6(f), we present attack comparisons with baselines evaluated on the CIFAR-10 and CIFAR-100 datasets. It is evident that BadSFL outperforms the baseline attacks in terms of both effectiveness and durability. To be specific, BadSFL achieves above 80% backdoor task accuracy across all types of backdoor attacks within the first 10 rounds, while the attacker remains active in the training process for backdoor training and malicious updates to the server. Meanwhile, BadSFL keeps the primary task accuracy at 60% (Figure 5(b)). Furthermore, even after the attacker exits the training process at the 40th round, the benign clients continue submitting normal updates in subsequent rounds, which could potentially affect the poisoned updates from the attacker in previous attacking rounds thus erasing the backdoor function. Despite this, BadSFL ensures a resilient backdoor function with accuracy exceeding 90% over the entire 100 SFL rounds, which is 3 times longer than the lifespan achieved by the two baseline attacks, where backdoor task accuracy drops below 50% after the 60th round. Horizontally comparing the effects of different types of backdoor trigger injections (Figures 6(a), 6(b) and 6(c)), it is observed the feature-based trigger performs the best among them, benefiting from its stealthiness without directly manipulating the images, thereby making its updates less likely to conflict with those from benign updates. We also provide more results available in Figure 7.
(a) Label Flipping, CIFAR-10, differential privacy + model pruning.
(b) Pattern Trigger, CIFAR-10, differential privacy + model pruning.
(c) Feature-based Trigger, CIFAR-10, differential privacy + model pruning.
(d) Feature-based Trigger, CIFAR-100, differential privacy + model pruning.
(e) Feature-based Trigger, CIFAR-100, FLAME.
(f) Feature-based Trigger, CIFAR-100, SparseFED.
@@ 5.3 Effectiveness of BadSFL
Figures 6(g) and 6(h) showcase the results obtained on the MNIST dataset. Similarly, BadSFL outperforms the other baseline attacks, achieving both backdoor task accuracy and primary task accuracy above 85%. After the malicious update injection stops at round 40, in the Label flipping attack, the backdoor task accuracy in two baseline attacks catastrophically drops below 40% within 10 rounds whereas BadSFL maintains a 5 times longer-lasting backdoor function in the global model in future rounds. In the pattern trigger attack, BadSFL also injects a more effective backdoor function into the global model with 10% higher accuracy compared to the baselines.
Figure 9: Neurotoxin.
Defense. We further conduct multiple defense strategies to demonstrate the superiority of BadSFL. Specifically, we use differential privacy [41], model pruning [24], FLAME [30] and SparseFed [32] to resist backdoor attacks. Figure 8 shows that BadSFL can still maintain high effectiveness under multiple defense strategies. Taking the label flipping on CIFAR-10 as an example, it can be observed that our method persistently maintains a high attack success rate even under defense strategies, while other baselines fail.
@@
However, our results do not confirm Neurotoxin’s expected effectiveness. Despite using this strategy, backdoor accuracy declines similarly to the baseline attack after the attacker exits at round 50. This trend is also observed in BadSFL experiments, warranting further investigation.
6 Conclusion